Find the mass of
a. the water molecules H2O and
b. The rthyl acohol molecules C2H6O. The atomic masses of H,. C, and O are respectively 1,808, 12.01 and 16.00U
1
Expert's answer
2019-06-06T10:05:02-0400

Solution.

m(H) = 1.808

m(C) = 12.01

m(O) = 16.00


"M(H2O) = 2 \times m(H) + 1 \times m(O)"

M(H2O) = 19.616


"M(C2H6O) = 2 \times m(C) + 6 \times m(H) + 1 \times m(O)"

M(C2H6O) = 50.868

Answer:

M(H2O) = 19.616

M(C2H6O) = 50.868